10. If 17th term of an A.P. exceeds its 10th term by 7. Find the common difference.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
We know that, for an A.P series; an = a+(n−1)d a17 = a+(17−1)d a17 = a +16d In the same way, a10 = a+9d As it is given in the question, a17 − a10 = 7 Therefore, (a +16d)−(a+9d) = 7 7d = 7 d = 1 Therefore, the common difference is 1.